User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

blog: katherine_maintenance_trip_june_15-24, 2014

Ke maintanance trip



- mk5's are supplied with new Smart power supplies: mk5ke with 1200W, mk5-2ke with 850W. Originally they were 550W units.

- mk5ke and mk5-2ke were successfully tested for recording and conditioning with two 16-TB modules.

- mk5ke and mk5-2ke: recording, conditioning, autocorrelation, bit stats and pcal tests (pcalmon) were performed with both mk5's, swapping VSI cables and FILA OUT VSI outputs.

- Pcal 5-MHz filter: tested, found to create fast phase wrap and introduce spurious signal in channels 5, 15, 16, thus discarded (pic – pcalmon output graphs). This was the crystal filter that Eric tested and improved.

- Pcal: Searched for spurious signals at image frequencies in S band, nothing found.

- Pcal: pcal 5-MHz signal comes now directly from the maser (5MHz - 2 connector) through 12-dB attenuator, bypassing the 5-MHz distributor. Levels regulated at the feed: added 3-dB attenuator to S-band, current attenuation at the feed is 15dB in S-band, 0dB in X band. Amplitudes of the pcal signal before DBBC: X-band ~18dB, S-band ~15dB, with spec an set to 10 KHz span, sample count 10 in averaging. 350 MHz for S, 1.1 GHz for X.

- DBBC: FILA OUT and Core2 boards inspected, connections checked. Secured VSI cables inside the DBBC with cable holders in order them not to pull on the VSI connectors.

Problems revealed:

  1. mk5ke: bank B back plane failure. Affected disks 4/5 and 6/7. Do not use for conditioning.
  2. mk5-2ke: module doesn't go all the way in to bank B.
  3. DBBC: “stuck bit” and bit statistics problems related to VSI OUT connectors: ch15 (S) with VSI1 (rear) – was problematic from the beginning, ch5 (X) with VSI2 (front) emerged after the inspection and refitting of FILA OUT board.

- DBBC: suppresses the amplitude of 10-kHz pcal tone (first) in each band.

- Maser: replaced the 'switches' unit and AC/DC converter, cavity tuning and frequency adjustment performed, new standby batteries were installed.


- mk5-ke is a main recorder connected to VSI2 on FILA OUT. Channel 5 have a stuck bit. - 5MHz signal for pcal is connected directly to the second output of the maser through the synch port.

Front end


- Antenna: Replaced jack screw elevation boot with zip-up type. Re-cut weather shield panels on elevation drive room to accommodate slightly larger zip-up boot.

- Antenna: attempted to remove elevation jack screw 2 1/2 inch pin to dish. Unsuccessful even with 10 ton hydraulic ram. Bronze bushing began to press out of housing following small movement of pin. May be sign that pin has become partially seized in bushing, however bushing does not appear to rotate with pin. Re-greased pin at nipples.

- Antenna: Removed elevation motor at the same time. Had to put nuts between motor and gearbox mating surfaces combined with mounting bolts bearing on mating surface to effect puller action. Extremely hard to pull. Anti-seize compound on shaft and key had turned to solid glue. Cleaned and refitted.

- Feed: Pcal levels regulated at the feed: added 3-dB attenuator to S-band, current attenuation at the feed is 15dB in S-band, 0dB in X band. - Feed: temperature sensor was installed and needs testing.


- Control room: replaced rotten door on control room, added outside seal at bottom to stop dust and water. New set of shelves in control room for modules.

- Control room: mk5 light for reading module labels is no longer connected to exterior light switch on ipske. It now switches on and is labelled light int

- Installed 2 of power distribution units (PDU). These are like an 8 way power board that have individually switched outputs. They can be switched from a web interface at the following 2 links:

The user is admin and the password is the same one we use for the internet power switches.
PDU1KE so far has wifi, cduke (network maint computer) and the interior light.
PDU2KE so far has mk5-2ke and the GPIB controller.

You could leave a comment if you were logged in.
/home/www/auscope/opswiki/data/pages/blog/blogkatherine_maintanance_trip_june_15-24.txt · Last modified: 2014/06/27 06:09 by Brett Reid